OpenStudy (anonymous):

inverse function of G(x) = -1/√ x

OpenStudy (anonymous):

y=-1/root(x). I will swap x and y round and then rearrange to find y.(my method to find inverse) x=-1/root(y) therefore root(y) = -1/x. square both sides to get y=1/x^2 so inverse of G is 1/x^2